obtain the riccati equation associated with the equation y''+w2y=0 and hence find its solution.

w=omega
1
Expert's answer
2019-03-28T10:55:35-0400

First, we consider some of the equalities associated with taking derivative of composite functions:


"y=y(x)\rightarrow
\left\lbrace\begin{array}{c}
\displaystyle\frac{d}{dx}\left(y^2\right)=2y\cdot y'\\[0.5cm]
\displaystyle\frac{d}{dx}\left(y'^2\right)=2y'\cdot y''
\end{array}\right."

Then,


"2y'\cdot\left|y''+\omega^2y=0\right.\rightarrow 2y'\cdot y''+\omega^2\cdot(2y'\cdot y)=0\rightarrow""\frac{d}{dx}\left(y'^2\right)+\frac{d}{dx}\left(\omega^2\cdot y^2\right)=0\rightarrow\frac{d}{dx}\left(y'^2+\omega^2\cdot y^2\right)=0\rightarrow""y'^2+\omega^2\cdot y^2=C^2\rightarrow\frac{dy}{dx}=\pm\sqrt{C^2-\omega^2\cdot y^2}\rightarrow""\frac{dy}{\sqrt{C^2-\omega^2\cdot y^2}}=\pm dx\rightarrow
\int\frac{dy}{C\cdot\sqrt{1-\left(\frac{\omega\cdot y}{C}\right)^2}}=\pm\int 1dx\rightarrow""\textit{We introduce a substitution}\\
u=\frac{\omega\cdot y}{C}\rightarrow du=\frac{\omega\cdot dy}{C}\rightarrow dy=\frac{C\cdot du}{\omega}\rightarrow"


"\int\frac{C\cdot du}{\omega\cdot C\cdot\sqrt{1-u^2}}=\pm x+C_1\rightarrow\frac{1}{\omega}\cdot\arcsin(u)=\pm x+C_1\rightarrow"

"\arcsin\left(\frac{\omega\cdot y}{C}\right)=\omega(\pm x+C_1)\rightarrow\frac{\omega\cdot y}{C}=\sin\left(C_1\omega\pm\omega x\right)\rightarrow"

"y(x)=\frac{C}{\omega}\cdot\sin\left(\overbrace{C_1\omega}^{A_1}\pm\omega\cdot x\right)"

Conclusion,


"\boxed{y(x)=\frac{C}{\omega}\cdot\sin\left(A_1\pm\omega\cdot x\right)}"

ANSWER

"y(x)=\frac{C}{\omega}\cdot\sin\left(A_1\pm\omega\cdot x\right)"